N₂ (g) + O₂ (g)<--> 2 NO (g)for which the K at 2273 K is 1.2 X 10⁻⁴. What is the concentration of NO if N₂=.166 M and O₂ =.145 M
1
Expert's answer
2020-06-08T15:15:14-0400

K = [NO]2/([N2]x[O2]) "\implies" [NO] = (Kx[N2]x[O2])1/2 = (1.2x10-4x0.166x0.145)1/2 = 0.0017 M